OpenStudy (yinshy):

What is the eighth term in the addition pattern that begins 13, 25, 37, 49? A. 85 B. 97 C. 109 D. 121

OpenStudy (yinshy):

@ijlal

OpenStudy (dayakar):

find the difference any two consecutive terms a2-a1 =25 -13 = 12 find a3-a2 =

OpenStudy (ijlal):

if you see the sequence 13,23,37,49 the difference is 12 so

OpenStudy (ijlal):